# Configurable photonic simulator for quantum field dynamics

### v1

@@ V.1 Quench dynamics
We consider the experimentally friendly vacuum state |$\ket{0}$ as the input state, in which case the OTA reduces to processing squeezed vacuum inputs with linear optics, as shown in Fig. 3. The input state describes zero photons, which is the ground state of the pre-quench Hamiltonian
@@
Importantly, |$\ket{0}$ belongs to the class of Gaussian states, which are characterized by their mean and covariance
@@
We time-evolve the state under the relativistic theory (30), which constitutes a quantum quench of both the mass and the couplings at $t=0$. We remark that the often-considered mass quench would instead require the input state to be the ground state of the Hamiltonian (30) with a different mass – thus describing no quasi-particles (with respect to the initial theory) instead. In the relativistic case, this would necessitate additional squeezers during state preparation, see e.g., Ref. [59].
During time evolution, the state remains Gaussian, with $r(t)=S(t)r_{0}=0$ and $\sigma(t)=S(t)\sigma S(t)^{\intercal}$ [84]. This greatly simplifies the study of quantum entanglement and correlations, as standard quantum information measures can be expressed in terms of the covariance matrix only. From an experimental point of view, the covariance matrix can be reconstructed through homodyne measurements [121], with the number of measurements being proportional to (sub-)system size. Multimode homodyne measurements, also relevant in the context of continuous-variable quantum computing and cluster state generation [122], have been demonstrated on various experimental platforms [123, 124, 125, 55, 54].
@@ V.2.1 Entanglement generation
We consider a spatial subregion $[0,\ell]$ with $\ell\leq L$, see inset of Fig. 4(a). As the modes are coupled, the degrees of freedom in $[0,\ell]$ become entangled with the ones in the complementary region $(\ell,L]$. Thus, while the global state $\bm{\rho}(t)$ remains pure during the unitary evolution, the local state $\bm{\rho}(t,l)=\textbf{Tr}_{L-\ell}\{\bm{\rho}(t)\}$ follows the evolution of an open system described by a completely positive trace-preserving map [126]. We assess entanglement generation via the Rényi-2 entanglement entropy, which, for Gaussian states, attains the simple form
@@
Figure 4: (a) Time evolution of the Rényi-2 entanglement entropy (47) (scaled by $10^{2}$) of the subregion $[0,\ell]$ following a quench from the optical vacuum (45) to the relativistic theory (30) with $\ell=L/5,m=1,\epsilon=2$ for varying $L=N\epsilon$ against the field-theoretic prediction (48) (black dashed curve). The light-cone-like motion of quasi-particles generates entanglement (see inset), leading to a linear increase in entanglement entropy until the time $\tau/N=\ell/(2Nv_{\text{max}})\approx 0.48$ (gray box) that the fastest quasi-particle pairs require to end up in distinct regions. (b) Same analysis for the Rényi-2 mutual information (49) between two spatial regions separated by a distance $d=L/5$. Correlations can only emerge after $t\geq\tau$, i.e., when the fastest quasi-particles have bridged the distance $d$.
Our main interest is to benchmark our quantum field simulator against QFT predictions. In the field theory context, the non-equilibrium dynamics of integrable models is understood in terms of a quasi-particle description [77, 78, 79, 80, 81, 82, 83] [see inset in Fig. 4a)]. Within this framework, the input state at $t=0$ acts as a source of quasi-particle pairs – momentum modes of opposite momenta $p$ and $-p$ – which ballistically spread correlations throughout the system after the quench. Then, at any given time $t>0$, a pair contributes to the entanglement entropy if one particle ends up in the subregion of interest, while the other resides in the complementary region. Combining this with the fact that the non-equilibrium steady-state for $t\to\infty$ attains the form of a generalized Gibbs ensemble (GGE) locally, i.e., $\bm{\rho}(\infty,l)\propto\text{exp}(-\sum_{p}y_{p}\bm{n}_{p})$, where $\bm{n}_{p}=\tilde{\bm{a}}^{\dagger}_{p}\tilde{\bm{a}}_{p}$ counts (conserved) quasi-particles with momenta $p$ and $y_{p}$ are Lagrange multipliers [130, 131], leads to the QFT prediction in the infinite volume limit
@@
Therein, $v(p)=\mathrm{d}\omega(p)/\mathrm{d}p=\sin(\epsilon p)/(\epsilon\omega(p))$ denotes the group velocity of the bosonic excitations and $s_{2}(p)$ is the thermodynamic Rényi-2 entropy density associated with the GGE. We refer to App. D.1 for finite-size corrections, to App. D.2 for an analytic calculation of $s_{2}(p)$, and to App. D.3 for a comment on divergences appearing in the continuum limit $\epsilon\to 0$.
@@
We compare the photonic simulation of the quench dynamics with the QFT prediction (48) (black dashed curve) in Fig. 4(a) for $\ell=L/5,m=1,\epsilon=2$ and varying system sizes $N=5-25$ (see App. B for circuit parameters when $N=5$). Shortly after the quench, we observe a linear growth of the Rényi-$2$ entanglement entropy $S_{2}(t,l)\propto t$ up to $t\leq\tau=\ell/(2v_{\text{max}})$, see gray region. The time $\tau$ marks the moment where two quasi-particles moving towards each other with the maximum velocity $v_{\text{max}}$ allowed by the relativistic dispersion have traveled a distance $\ell$ (see overlapping light cones in the inset).
@@
After $t\geq\tau$, also the slower quasi-particles start to contribute to the entanglement between the subregion $[0,\ell]$ and its complement, but the growth of the entanglement entropy slows down due to saturation effects. For an infinitely extended system, the entropy would settle to its stationary value dictated by the GGE. For a finite system with periodic boundaries, however, the quasi-particles return to their initial positions every time they have traversed the system. This leads to so-called revivals in the entanglement entropy, marked by a temporary dip around the time $L/(2v_{\text{max}})=5\tau$ the fastest quasi-particles took to cross the system once. As the quasi-particles’ momenta vary, the entanglement is never entirely destroyed, and the dip attains a finite width.
@@
Over the whole time range considered, the photonic simulation is in remarkable agreement with the QFT prediction – even for small system sizes $N\sim 10$ [51, 54]. Both the initial linear scaling and the revival dip(s) are particularly robust features that are clearly visible for even smaller mode numbers. Importantly, the circuit parameters fall into experimentally accessible regimes, see App. B. In general, implementing strong squeezing is most challenging. Here, the maximum squeezing required $z_{\text{max}}$, which corresponds to the zero-momentum mode, is independent of system size and evaluates to $\lvert z_{\text{max}}\rvert\approx 0.5\approx 4.3$dB – far below current records around $15$dB [132]. This demonstrates that the photonic simulation of entanglement dynamics in quantum fields is well within the reach of current experimental technologies.
@@ V.2.2 Correlation spreading
where $S_{2}(t,\ell\cup\ell)$ and $\sigma(t,\ell\cup\ell)$ are associated with the combined region $[0,\ell]\cup[\ell+d,2\ell+d]$. Analogously to the Rényi-$2$ entanglement entropy, the latter equals the Wigner mutual information for Gaussian states [127, 128, 129].
@@
Strikingly, the early period $t\lesssim 3\tau$ is well captured by simulations with small system sizes $N\sim 5-10$. We find deviations from the QFT prediction around the revival time. Here, reasonable agreement requires around $N\sim 20$ modes, which is still accessible on existing platforms.
@@ V.3 Systematic study: Bending of light cones
Algebraic $t\gtrsim d^{\gamma},\,\gamma\in(0,1)$ $21\leq\alpha<$ 2
Logarithmic $t\gtrsim\log d$ $10<\alpha<$ 1
@@
In the regime $21\leq\alpha<$ 2, see Figs. 5(c)-(d), the light cones start to bend, demonstrating faster correlation propagation for large distances, as expected for long-range couplings characterized by $f_{M}>0$ for all $M$. More precisely, the light cones follow algebraic curves. Simultaneously, nearest-neighbor correlations appear delayed, which is caused by a decreasing coupling strength between neighboring modes, as shown in the upper curve of the inset in Fig. 5(a). Upon further increasing the coupling range to $10<\alpha<$ 1, the light cones become logarithmic, see Fig. 5(e), until they disappear in the limit $\alpha\to 0$, see Fig. 5(f). The latter case corresponds to an all-to-all coupling of the modes with $f_{M}\to 1/2$ for all distances $M$, allowing for immediate information transfer and rendering all modes pari passu.
@@
VI Discussion

### v2

@@ V.1 Quench dynamics
[…]
@@
[…]
@@
We time-evolve the state under the relativistic theory (31), which constitutes a quantum quench of both the mass and the couplings at $t=0$. We remark that the often-considered mass quench would instead require the input state to be the ground state of the Hamiltonian (31) with a different mass—thus describing no quasi-particles (with respect to the initial theory) instead. In the relativistic case, this would necessitate additional squeezers during state preparation, see e.g., Ref. [99].
During time evolution, the state remains Gaussian, with $r(t)=S(t)r_{0}=0$ and $\sigma(t)=S(t)\sigma S^{\intercal}(t)$ [159]. This greatly simplifies the study of quantum entanglement and correlations, as standard quantum information measures can be expressed in terms of the covariance matrix only. From an experimental point of view, the covariance matrix can be reconstructed through homodyne measurements [48], with the number of measurements being proportional to (sub-)system size. Multimode homodyne measurements, also relevant in the context of continuous-variable quantum computing and cluster state generation [57], have been demonstrated on various experimental platforms [11, 76, 33, 82, 5].
@@ V.2.1 Entanglement generation
We consider a spatial subregion $0,\ell]$ with $\ell\leq L$, see inset of Fig. 4(a). As the modes are coupled, the degrees of freedom in $0,\ell]$ become entangled with the ones in the complementary region $(\ell,L]$. Thus, while the global state $\bm{\rho}(t)$ remains pure during the unitary evolution, the local state $\bm{\rho}(t,l)=\textbf{Tr}_{L-\ell}\{\bm{\rho}(t)\}$ follows the evolution of an open system described by a completely positive trace-preserving map [116]. We assess entanglement generation via the Rényi-2 entanglement entropy, which, for Gaussian states, attains the simple form
@@
Our main interest is to benchmark our quantum field simulator against QFT predictions. In the field theory context, the non-equilibrium dynamics of integrable models is understood in terms of a quasi-particle description [34, 35, 36, 37, 6, 38, 103] [see inset in Fig. 4(a)]. Within this framework, the input state at $t=0$ acts as a source of quasi-particle pairs—momentum modes of opposite momenta $p$ and p-p —which ballistically spread correlations throughout the system after the quench. Then, at any given time $t>0$, a pair contributes to the entanglement entropy if one particle ends up in the subregion of interest, while the other resides in the complementary region. Combining this with the fact that the non-equilibrium steady-state for $t\to\infty$ attains the form of a generalized Gibbs ensemble (GGE) locally, i.e., $\bm{\rho}(\infty,l)\propto\text{exp}(-\sum_{p}y_{p}\bm{n}_{p})$, where $\bm{n}_{p}=\tilde{\bm{a}}^{\dagger}_{p}\tilde{\bm{a}}_{p}$ counts (conserved) quasi-particles with momenta $p$ and $y_{p}$ are Lagrange multipliers [93, 152], leads to the QFT prediction in the infinite volume limit
@@
Therein, $v(p)=\mathrm{d}\omega(p)/\mathrm{d}p=\sin(\epsilon p)/(\epsilon\omega(p))$ denotes the group velocity of the bosonic excitations and $s_{2}(p)$ is the thermodynamic Rényi-2 entropy density associated with the GGE. We refer to App. D.1 for finite-size corrections for periodic and open boundaries, to App. D.2 for an analytic calculation of $s_{2}(p)$, and to App. D.3 for a comment on divergences appearing in the continuum limit $\epsilon\to 0$.
Figure 4: (a) Time evolution of the Rényi-2 entanglement entropy (51) (scaled by $10^{2}$) of the subregion $0,\ell]$ following a quench from the optical vacuum (49) to the relativistic theory (31) with $\ell=L/5,m=1,\epsilon=2$ for varying $L=N\epsilon$ against the field-theoretic prediction (52) (black dashed curve). The light-cone-like motion of quasi-particles generates entanglement (see inset), leading to a linear increase in entanglement entropy until the time $\tau/N=\ell/(2Nv_{\text{max}})\approx 0.48$ (gray box) that the fastest quasi-particle pairs require to end up in distinct regions. (b) Same analysis for the Rényi-2 mutual information (53) between two spatial regions separated by a distance $d=\ell$. Correlations can only emerge after $t\geq\tau$, i.e., when the fastest quasi-particles have bridged the distance $d$. We refer to Fig. 10 in App. D.1.2 for a treatise of open boundaries.
@@
We compare the photonic simulation of the quench dynamics with the QFT prediction (52) (black dashed curve) in Fig. 4(a) for $\ell=L/5,m=1,\epsilon=2$ and varying system sizes N=525N=5-25 (see App. B for circuit parameters when $N=5$). Shortly after the quench, we observe a linear growth of the Rényi-$2$ entanglement entropy $S_{2}(t,l)\propto t$ up to $t\leq\tau=\ell/(2v_{\text{max}})$, see gray region. The time $\tau$ marks the moment where two quasi-particles moving towards each other with the maximum velocity $v_{\text{max}}$ allowed by the relativistic dispersion have traveled a distance $\ell$ (see overlapping light cones in the inset).
@@
After $t\geq\tau$, also the slower quasi-particles start to contribute to the entanglement between the subregion $0,\ell]$ and its complement, but the growth of the entanglement entropy slows down due to saturation effects. For an infinitely extended system, the entropy would settle to its stationary value dictated by the GGE. For a finite system with periodic boundaries, however, the quasi-particles return to their initial positions every time they have traversed the system. This leads to so-called revivals in the entanglement entropy, marked by a temporary dip around the time $L/(2v_{\text{max}})=5\tau$ the fastest quasi-particles took to cross half of the system. As the quasi-particles’ momenta vary, the entanglement is never entirely destroyed, and the dip attains a finite width.
@@
Over the whole time range considered, the photonic simulation is in remarkable agreement with the QFT prediction—even for small system sizes $N\sim 10$ [98, 5]. Both the initial linear scaling and the revival dip(s) are particularly robust features that are clearly visible for even smaller mode numbers. This also holds true for open boundary conditions, in which case circulanticity is broken, see App. D.1.2. Importantly, the circuit parameters fall into experimentally accessible regimes, see App. B. In general, implementing strong squeezing is most challenging. Here, the maximum squeezing required $z_{\text{max}}$, which corresponds to the zero-momentum mode, is independent of system size and evaluates to $\lvert z_{\text{max}}\rvert\approx 0.5\approx 4.3$dB—far below current records around 1515dB [151]. This demonstrates that photonic simulations of entanglement dynamics in quantum fields are well within the reach of current experimental technologies.
@@ V.2.2 Correlation spreading
where $S_{2}(t,\ell\cup\ell)$ and $\sigma(t,\ell\cup\ell)$ are associated with the combined region $0,\ell]\cup[\ell+d,2\ell+d]$. Analogously to the Rényi-$2$ entanglement entropy, the latter equals the Wigner mutual information for Gaussian states [71, 52, 53].
@@
Strikingly, the early period $t\lesssim 3\tau$ is well captured by simulations with small system sizes $N\sim 5-10$. We find deviations from the QFT prediction around the revival time. Here, reasonable agreement requires around $N\sim 20$ modes, which is still accessible on existing platforms. Both findings extend to open boundaries, see App. D.1.2.
@@ V.3 Systematic study: Bending of light cones
Algebraic $t\gtrsim d^{\gamma},\,\gamma\in(0,1)$ $1\leq\alpha<2$
Logarithmic $t\gtrsim\ln d$ $0<\alpha<1$
@@
In the regime $1\leq\alpha<2$, see Figs. 5(c)-(d), the light cones start to bend, demonstrating faster correlation propagation for large distances, as expected for long-range couplings characterized by $f_{M}>0$ for all $M$. More precisely, the light cones follow algebraic curves. Simultaneously, nearest-neighbor correlations appear delayed, which is caused by a decreasing coupling strength between neighboring modes, as shown in the upper curve of the inset in Fig. 5(a). Upon further increasing the coupling range to $0<\alpha<1$, the light cones become logarithmic, see Fig. 5(e), until they disappear in the limit $\alpha\to 0$, see Fig. 5(f). The latter case corresponds to an all-to-all coupling of the modes with $f_{M}\to 1/2$ for all distances $M$, allowing for immediate information transfer and rendering all modes pari passu.
@@
VI Error analysis
Photonic devices, like any experimental realization, have imperfections. Therefore, it is crucial to study the effects of experimental errors on our benchmarks. We focus on the two predominant error mechanisms in Boson-Sampler setups (see Fig. 3): noise due to parameter fluctuations during state preparation and photon loss within the interferometer [136]. We investigate the relativistic theory considered in Sec. V.2, see Fig. 4.
VI.1 State preparation noise
The state entering the interferometer is a squeezed vacuum state with a phase. During its preparation, both the intensity $\xi$ and the phase $\Phi$ fluctuate from one experimental run to another. We remark that numerical simulations with tensor networks indicate that Boson Sampling remains classically hard in the presence of phase fluctuations [115]; see also [140] for a recent sampling experiment discussing phase noise.
We simulate noise by randomly selecting the optical parameters from a Gaussian distribution centered around the ideal values. Figure 6 shows the results obtained with $\mathcal{N}_{\text{S}}=100$ such noise realizations. The standard deviation $\Delta$ of the distribution determines the noise strength. As we expect fluctuations in the squeezing amplitude to increase with the amount of squeezing, we consider $\Delta_{\xi}=\xi\delta$ (left column). In contrast, we sample the phase from a Gaussian with constant width $\Delta_{\Phi}=\delta$ (right column). In both cases, $\delta>0$ is a noise parameter with $\delta=0$ corresponding to an ideal, noiseless input.
We distinguish two scenarios characterized by the time scales over which parameters fluctuate. If the input state can be prepared repeatedly with stable experimental parameters over the measurement time scale for a given set of model parameters, e.g., one time step of the evolution, we refer to the noise as low frequency, see Figs. 6(a)-(d). In this case, we directly compute information measures for each of the $\mathcal{N}_{\text{S}}$ randomly drawn parameters and average the result, e.g., $\langle S_{2}(\sigma)\rangle$. In a second scenario, circuit parameters fluctuate on time scales shorter than the measurement window, thereby affecting the estimation of individual covariance matrices and thus referred to as high-frequency, see Figs. 6(e)-(h). Here, we instead average the covariance $\sigma$ over $\mathcal{N}_{\text{S}}$ noise realizations before computing the quantities of our interest, e.g., $S_{2}(\langle\sigma\rangle)$. We note that low-frequency noise can always be converted to high-frequency noise by randomizing the ordering of covariance matrix measurements and model parameter choices.
Figure 6: Effects of state preparation noise for the relativistic quench, cf. Sec. V.2. We consider the same setup underlying Fig. 4, that is, $N=25$ modes, a subinterval of length $\ell=L/5$ for the Rényi-2 entropy and two intervals of the same length separated by a distance $d=\ell$ for the Rényi-2 MI, as well as $m=1$, $\epsilon=2$. We sample $\mathcal{N}_{\text{S}}=100$ realizations of the normally-distributed optical parameters $\xi$ (squeezing amplitude, left column) and $\Phi$ (phase, right column) with fluctuations controlled via the noise parameter $\delta\in\{0,\,0.1,\,0.2,\,0.3\}$. The shaded areas cover one standard deviation. For high-frequency noise, we resampled $100$ times to obtain the $\pm 1$ s.d. interval, which turned out smaller than the line width. All information measures shift upwards for low-frequency noise, see panels (a)(d), while high-frequency noise mimics mixing, which increases local entropy and decreases correlations, as shown in panels (e)(h). All qualitative features of quench dynamics are preserved up to noise levels of $30\%$.
For low-frequency noise, we find that both the entropy and the mutual information slightly increase with respect to the noiseless case. This can be understood by considering the relation between the number of quasi-particles present, $n$, and the optical parameters. For the relativistic theory (30), a momentum mode $k$ is populated by $n_{k}=\sinh^{2}z_{k}$ quasi-particles, see Eq. (101), discussed in App. D.2, with Eq. (16b) and $\gamma_{k}=\epsilon\omega_{k}$ understood. Hence, symmetrically distributed deviations from the ideal values for $\xi$ and $\Phi$ are biased towards higher quasi-particle numbers, resulting in an overall growth of correlations.
For high-frequency noise, the observed increase in entropy and decrease in mutual information is explained by their elementary properties: mixing probability distributions increases their entropy and washes out correlations [111]. Remarkably, the mutual information remains almost unaffected by amplitude noise, see Figs. 6(g). In fact, the covariance matrix of a (complex) squeezed vacuum state has diagonal elements $\cosh(2\xi)\mp\cos(\Phi)\sinh(2\xi)$ and off-diagonal elements $\sin(\Phi)\sinh(2\xi)$. Hence, averaging over different amplitudes affects the covariance matrix elements in similar ways. When inserting the latter into the defining equation (53), this causes cancellations of fluctuations to first order. In contrast, phase fluctuations do not correspond to global modifications.
VI.2 Photon Loss
Losses are the main bottleneck of photonic platforms [136]. Most photons are lost in the interferometer due to the many integrated beam splitters and phase shifts. Over the past decade, losses have posed significant challenges to the race toward quantum advantage, as sufficiently high loss levels can render boson sampling efficient on classical hardware [32, 113, 114, 43].
The standard loss model in quantum optics couples each mode emerging from an (ideal) beam splitter to an ancillary vacuum mode via a beam splitter with transmissivity $T\in[0,1]$, which controls the loss rate ($T=1$ amounts to the lossless case). This is referred to as the symmetric loss model, see the inset in Fig. 7(a) for a sketch. The corresponding quantum channel acts on every two-mode covariance matrix $\sigma$ as [134]
$\varepsilon(\sigma)=TS_{\text{BS}}\sigma S_{\text{BS}}^{\intercal}+\tfrac{1}{2}(1-T)\mathds{1}_{4}.$ (55)
For a generic input covariance matrix $\sigma$, we compute $\varepsilon(\sigma)$ using a Python package available under [107].
An important factor influencing the impact of losses on the simulation is the interferometer geometry. Given the structural differences between the two predominant schemes, namely Reck’s triangular [125] and Clement’s [45] square meshes, we study both cases. For the former, we further distinguish between configurations where the beam splitters are arranged as a triangle pointing upwards and those where it points downwards (see Fig. 8 in App. A.2.3 for a pictorial representation of the down-oriented configuration), while the latter is homogeneous by construction.
In Fig. 7, we show the results for the relativistic quench and the three interferometer configurations (Reck up: blue triangles, Clements: green squares, Reck down: yellow triangles) at $2\%$ losses per mode, which translates into an overall loss of $\sim 38.4\%$ for the Clements decomposition. This corresponds to typical loss rates of commercially available multi-mode interferometers [140]. For all three decompositions, the lossy data are in qualitative agreement with the ideal case (purple points). The Rényi entropy, see Fig. 7(a), remains robust to loss, which we attribute to two competing effects. On the one hand, loss decreases overall state purity and thus increases entropy. On the other hand, losses decrease the mean photon number, rendering the states more similar to the vacuum, and thus decreasing subsystem entropy. In contrast, losses generally reduce correlations, as seen in Fig. 7(b).
Figure 7: Effects of a lossy interferometer for the relativistic quench, cf. Sec. V.2. We consider the same setup underlying Fig. 4, that is, $N=25$ modes, a subinterval of length $\ell=L/5$ for the Rényi-2 entropy, see (a), and two intervals of the same length separated by a distance $d=\ell$ for the Rényi-2 mutual information, see (b), as well as $m=1$, $\epsilon=2$. We apply a loss channel to both output modes of each beam splitter by interfering with the vacuum at another beam splitter with transmittivity $T=0.98$ (see inset), resulting in $\sim 61.6\%$ overall transmittance. We compare three decomposition geometries (Reck up: blue triangles, Clements: green squares, Reck down: yellow triangles) with the ideal case (purple points). While the entropy remains robust to losses, having fewer quasi-particles reduces the mutual information. Nevertheless, qualitative features remain unaffected by losses.
We further observe a consistent ordering of information measures with respect to the decomposition scheme: both quantities of our interest are largest (smallest) for the upward (downward) Reck scheme, while Clements’ scheme is positioned in between. This is because we consider the entropy of the 5-mode subsystem formed by the first five modes located at the upper end of the system. Hence, when the tip of the Reck triangle points upward, less noisy beam splitters are encountered. Instead, Clements’ decomposition does not break spatial translation symmetry, and thus loss affects all modes equally. Homogeneous losses make loss effects easier to interpret and control, which is why Clements’ decomposition is often preferred.
In summary, none of the considered experimental error sources alter the qualitative features of the simulated information dynamics, and error-induced deviations are moderate in experimentally realistic parameter regimes.
VII Discussion